OpenStudy (anonymous):

(x/5)-2=(x/5) solve for x

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

Si!

OpenStudy (anonymous):

hi can you help me solve this problem 2(x+3)+x=4(x-3)

OpenStudy (anonymous):

2x + 6 +x = 4x -12 3x +6 = 4x-12 -x = -18 x=18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

next question 160=-50x-40